Alex has 100 friends on Snapchat his goal is to add five new friends a day a) write an expression that represents his total number of friends make sure to define your variable b) solve your expression for the total number of friends he will have after 10 days


1
Expert's answer
2021-11-03T17:08:39-0400

a) Let "x=" the number of days, "y=" total number of friends. Then an expression that represents his total number of friends is


"y(x)=100+5x"

b)


"y(10)=100+5(10)=150"

He will have 150 friends after 10 days.